# An Exploration of Default Images in Text-to-Image Generation

### v4

@@ 6.5. Limitations and Future Work
Our focus on Midjourney limits the external validity of findings to other TTI systems. We selected this scope to allow in-depth qualitative and quantitative analysis within a consistent model environment. Future research could adopt a comparative lens, investigating default images in systems such as Stable Diffusion (Rombach et al., 2022), Imagen (Saharia et al., 2022), Flux (Labs et al., 2025), Janus (Chen et al., 2025), or emerging video models, potentially revealing differences in model capabilities and biases beyond canonical benchmarks.
@@
Due to the vastness of Midjourney’s latent space, our analysis cannot be exhaustive (Quaye et al., 2025; Liu and Chilton, 2022a). Each model version introduces its own set of default images, with some occurring far more frequently than others (cf. Figure 6). As a result, our work can only provide a partial view of the phenomenon. Nonetheless, certain defaults (like the Lady-Birdhead image), are common and straightforward to identify even in small-scale manual studies, suggesting that foundational insights can be gained without full coverage. Future work could extend our approach to systematically map default images across different model versions. Default images could form the basis of a benchmark for model evaluation, capturing how systems respond to unknown or ambiguous inputs. Such a benchmark would complement existing technical metrics by foregrounding model behavior in edge-case scenarios encountered in creative practice.
@@
We restricted our analysis to short prompts to maximize the likelihood of encountering default images while keeping computational demands manageable. This constraint limits conclusions about default images that may emerge in more elaborate prompt contexts. Future studies could explore how prompt length, structure, modifiers, and semantic specificity influence the emergence and variability of default images. From a preliminary study, we saw different images arise when combining large amounts of our prompts. Future work could study whether these prompts are default images, or whether there is a change in model behavior when prompted with longer prompts. Negative prompts and weighted prompts were also not in scope of our foundational study and left to future work. Further, our sensitivity analysis and ablation studies were limited and offer new ground for research. Future work could systematically examine the role of seeds, negative prompts, and other parameters in shaping default outputs.
@@
We did not directly investigate bias in default images in this work, yet their occurrence in the absence of explicit text guidance potentially provides an opportunity to study “raw” biases in TTI models. Building on related studies on bias in TTI models (e.g., (Vasilev et al., 2024; Elsharif et al., 2025; Saravanan et al., 2023; Kandwal and Nehra, 2024; Luccioni et al., 2023; Lan et al., 2025; Jha et al., 2024; Vice et al., 2023)), future research could analyze demographic, cultural, or aesthetic biases embedded in default images.
Studying default images from a color theory perspective could also provide new insights on prevailing harmonies and contrasts and contribute to an understanding of the symbolism and potential impact of such associations. Exploring how default images tie into the creative practice could be an interesting avenue for future studies. Future work could examine the implications of default images for creativity support, for example through Construal Level Theory (Trope and Liberman, 2010) or other psychological frameworks.
@@ 7. Conclusion
We presented the first systematic investigation of default images in Midjourney –– images with a certain motif and style that occur in response to a variety of prompts. Through an initial study with manual prompt‐generation and affinity diagramming, we identified ten canonical default images. Scaling up to a larger set of images, we presented a simple method for identifying default images in Midjourney. Through a large-scale computational analysis of Midjourney images, we demonstrated that default images are not just a theoretical construct, but occur in response to real-world user prompts. Our user study showed that default images may negatively influence user satisfaction. Based on our findings, we distilled eight postulates characterizing the properties of default images. By defining the notion of default images, our work lays the groundwork for future research on the limits of text‐to‐image systems.

### v5

@@ 6.6. Limitations and Future Work
We presented a systematic empirical study without having access to the underlying training data or model. It is our belief that default images arise from latent model behavior, likely due to gaps in the training data. However, some system internals on Midjourney’s platform (like the sampling mechanism or potential prompt rewriting) may contribute to the phenomenon of default images. For instance, the many-to-many relationship between prompts and images observed in Figure 2 can be attributed to diversity mechanisms in Midjourney’s stochastic sampling. On the other hand, other mechanisms, such as the LLM-based moderation feature, can likely be excluded as an alternative explanation for default images. Moderation either lets prompts pass or not, and moderation is therefore not mechanistically involved in generating default images. Nevertheless, important interpretative limitations around causality remain due to the opaque nature of Midjourney’s proprietary platform. Without knowing whether and how the system is intervening in its inputs and outputs, it is difficult to fully interpret whether default images reflect model behavior or platform design decisions.
@@
Our focus on Midjourney limits the external validity of findings to other TTI systems. We selected this scope to allow in-depth qualitative and quantitative analysis within a consistent model environment. Future research could adopt a comparative lens, investigating default images in other TTI generation systems, such as Stable Diffusion (Rombach_2022_CVPR), Imagen (2205.11487.pdf), Flux (labs2025flux1kontextflowmatching), Janus (janus), or video generation models, potentially revealing differences in model capabilities and biases beyond canonical benchmarks.
@@
Due to the vastness of Midjourney’s latent space, our analysis cannot be exhaustive (2507.17922.pdf; 3491102.3501825.pdf). Each model version introduces its own set of default images, with some occurring far more frequently than others (cf. Figure 6). As a result, our work can only provide a partial view of the phenomenon. Nonetheless, certain defaults (like the Lady-Birdhead image), are common and straightforward to identify even in small-scale manual studies, suggesting that foundational insights can be gained without full coverage. Future work could extend our approach to systematically map default images across different model versions. Default images could form the basis of a benchmark for model evaluation, capturing how systems respond to unknown or ambiguous inputs. Such a benchmark would complement existing technical metrics by foregrounding model behavior in edge-case scenarios encountered in creative practice.
@@
We restricted our analysis to short prompts to maximize the likelihood of encountering default images while keeping computational demands manageable. This constraint limits conclusions about default images that may emerge in more elaborate prompt contexts. Future studies could explore how prompt length, structure, modifiers, and semantic specificity influence the emergence and variability of default images. From a preliminary study, we saw different images arise when combining large amounts of our prompts. Future work could study whether these prompts are default images, or whether there is a change in model behavior when prompted with longer prompts. Negative prompts, weighted prompts, and image prompts were also not in scope of our foundational study and left to future work. Further, our sensitivity analysis and ablation studies were limited and offer new ground for research. Future work could systematically examine the role of seeds, negative prompts, and other parameters in shaping default outputs.
@@
We did not directly investigate bias in default images in this work, yet their occurrence in the absence of explicit text guidance potentially provides an opportunity to study “raw” biases in TTI models. Building on related studies on bias in TTI models (e.g., (Vasilev2024S137; Elsharif2025122636; Saravanan202384; Kandwal202473; luccioni2023stablebiasanalyzingsocietal; 10.1145/3706599.3719678; jha2024visageglobalscaleanalysisvisual; vice2023quantifyingbiastexttoimagegenerative)), future research could analyze demographic, cultural, or aesthetic biases embedded in default images.
Studying default images from a color theory perspective could also provide new insights on prevailing harmonies and contrasts and contribute to an understanding of the symbolism and potential impact of such associations. Exploring how default images tie into the creative practice could be an interesting avenue for future studies. Future work could examine the implications of default images for creativity support, for example through Construal Level Theory (Trope2010-TROCTO) or other psychological frameworks.
@@ 7. Conclusion
We presented the first systematic investigation of default images in Midjourney. Default images are images with a certain motif that occur in response to a variety of dissimilar prompts. Through an initial study with manual prompt‐generation and affinity diagramming, we identified ten canonical default images. Scaling up to a larger set of images, we presented a simple method for identifying default images in Midjourney. Through a large-scale computational analysis of Midjourney images, we demonstrated that default images are not just a theoretical construct, but occur in response to real-world user prompts. Our user study showed that default images may negatively influence user satisfaction. Based on our findings, we distilled eight postulates characterizing the properties of default images. By defining the notion of default images, our work lays the groundwork for future research on the limits of generative systems.